OpenStudy (anonymous):

The coordinates of quadrilateral ABCD are A(-1,5),B(8,2),C(11,13),D(2,6). The midpoint of AC is .....,....... The midpoint of BD is ......,....... The slope of AC is ........., The slope of BD is ........., Is quadrilateral ABCD is a rhombus ?[Y,N]----.

OpenStudy (anonymous):

mid point of AC = (-1+11/2, 5+13/2) = (10/2, 18/2) = (5,9) Mid point of BD = (8+2/2, 2+6/2) = (10/2, 8/2) = (5,4) Slope of AC = 13-5/11+1 = 8/12 = 2/3. slope of BD = 6-2/2-8 = 4/-6 = -2/3 We know that rhombus has a common mid point of its diagonals and product of slopes = -1., not observed here. conclusion = ABCD not a rhombus

OpenStudy (anonymous):

many many thanks